# Common Inpainted Objects In-N-Out of Context

### v1

@@ Abstract
We present Common Inpainted Objects In-N-Out of Context (COinCO), a novel dataset addressing the scarcity of out-of-context examples in existing vision datasets. By systematically replacing objects in COCO images through diffusion-based inpainting, we create 97,722 unique images featuring both contextually coherent and inconsistent scenes, enabling effective context learning. Each inpainted object is meticulously verified and categorized as in- or out-of-context through a multimodal large language model assessment. Our analysis reveals significant patterns in semantic priors that influence inpainting success across object categories. We demonstrate three key tasks enabled by COinCO: (1) training context classifiers that effectively determine whether existing objects belong in their context; (2) a novel Objects-from-Context prediction task that determines which new objects naturally belong in given scenes at both instance and clique levels, and (3) context-enhanced fake detection on state-of-the-art methods without fine-tuning. COinCO provides a controlled testbed with contextual variations, establishing a foundation for advancing context-aware visual understanding in computer vision and image forensics. Our code and data are at: https://github.com/YangTianze009/COinCO.
@@ 1 Introduction
Context is fundamental to visual understanding [1, 2, 3]. When humans view a scene, we instinctively assess the contextual coherence between objects and their environment. This context-based reasoning is essential for interpreting real-world scenes [4] and beyond. Take a look at Figure 2—can you spot the inpainted (i.e., fake) object in each image? Contextual understanding helps identify objects that violate scene expectations, even when pixel-level artifacts are imperceptible. By using context, humans naturally assess whether objects appear in plausible settings—a potted plant flying in the sky would immediately raise suspicion, while the same plant in a garden appears perfectly natural.
@@
Learning context from data, however, is difficult. A significant challenge is that unusual scenes with out-of-context objects are rare in real life. Common computer vision datasets [5, 6, 7, 8] primarily contain objects in their natural settings, creating a scarcity of examples with contextual violations. This presents a fundamental obstacle: how can we train machine learning models that require large data to recognize contextual inconsistencies when such examples are inherently uncommon?
@@
To address this challenge, we propose a dataset with both in-context and out-of-context objects through controlled image manipulation. By systematically replacing objects in existing real scenes, we can generate the necessary examples of contextual violations while preserving the overall structure of the scene. We specifically chose the Common Objects in Context (COCO) dataset [5] as our foundation because it contains a diverse set of everyday scene photographs with rich image- and object-level annotations, offering an ideal starting point for contextual manipulation.
Figure 1: Which object is fake? Only one object per image is inpainted. Out-of-context inpainted objects are easier to identify. Answers are revealed at the bottom of this page [^2] Figure 2 answers. 1st row: bird, laptop; 2nd row: hotdog, potted plant, potted plant; 3rd row: horse, horse..
@@
Using Stable Diffusion’s inpainting model [9], we replace exactly one object per COCO image. This selective approach allows us to maintain the broader scene context while introducing precise, controlled variations in object-scene relationships. Our meticulous data creation pipeline and multi-step verification ensure high-quality inpainting results, and we apply a state-of-the-art Multimodal Large Language Model (MLLM) to classify each inpainted object as either contextually consistent or inconsistent with its scene.
@@
Figure 2 shows the pipeline of our data creation and the downstream tasks. We name our dataset Common Inpainted Objects In-N-Out of Context (COinCO), reflecting our emphasis on context for inpainted objects. Our dataset contains 97,722 unique inpainted images, each accompanied by rich annotations.
@@
COinCO is the first dataset that features inpainted objects with diverse context annotations, enabling advancements in tasks including context classification, object-from-context prediction, and fake detection (as our objects are inpainted by a generative model). COinCO serves as a challenging benchmark, encouraging new research for the computer vision community.
The main contributions of our work are:
@@
2. We employ COinCO’s context labels to train a context classifier. This task utilizes the visual features of an image, alongside the semantic features of detected objects to predict an object to be in- or out-of-context.
3. We introduce a novel task, object-from-context, which aims to predict instance- and clique-level categories for possible objects that fit the given context.
4. We demonstrate that context can be effectively integrated into fake detection pipelines, substantially improving state-of-the-art methods without fine-tuning.
@@ 2 Related Work
Context reasoning. Context is essential for understanding object relationships and finding anomalies in complex scenes. Biederman et al. [4] identified five relational principles—support, interposition, probability, position, and size—that help identify contextual inconsistencies. Prior works apply these principles using object co-occurrence and support relationships for out-of-context detection [2]. Others have demonstrated that slight context changes in “object transplanting” experiments can cause significant errors in object detection. Acharya et al. [15] introduced a Graph Contextual Reasoning Network (GCRN) to model co-occurrence and relative position to detect out-of-context objects. A comprehensive review of context in vision is in [16]. Ours is the first work to feature a mixture of in- and out-of-context objects in large-scale data. Moreover, context is under-explored in fake detection. Multimodal Large Language Models (MLLMs)  [17, 18, 19, 20, 21], offer a human-aligned approach to nuanced context reasoning. We are the first to use context in fake detection with MLLMs.
Figure 2: Our COinCO pipeline. (a) For a given COCO image, an object is randomly replaced by with Stable Diffusion inpainting. (b) Inpainting success is verified using object detection and an MLLM. Successes are added to the dataset, while fail cases are regenerated and retested. (c) Inpainted images are classified as in-context or out-of-context using the MLLM. (d) Instance-level (object category) and clique-level (supercategory) information is predicted from the inpainted image using an MLP and recovered with Stable Diffusion. (e) Fake pixels are detected with pretrained fake detectors and enhanced within the bounding box region of a detected out-of-context object. (f) Semantic and visual embeddings are derived from the object list and inpainted image, respectively. These embeddings are used to train an MLP for classifying instances as in- or out-of-context.
Fake image generation and detection. Image manipulation has a long history from traditional object insertion techniques [22, 23] to GAN-based synthesis [24, 25, 26, 27, 28]. Recently, diffusion models set new standards [29, 30], with text-to-image models like Stable Diffusion [9] and DALL-E [31] ControlNet enhances diffusion models with controls [32]. Emerging techniques tailor diffusion methods for object manipulation, inpainting, and harmonization [33, 34], 3D [35], and relighting [36, 37].
For detecting manipulated images, earlier works classify at image-level [38, 39]. Advanced methods detect various manipulations with details: PSCC-Net [40] employs spatio-channel correlation across scales, CAT-Net [41] focuses on JPEG compression artifacts, and ManTra-Net [42] captures diverse manipulation traces. TruFor combines RGB and noise-sensitive fingerprints to produce anomaly and confidence maps [43]. For fake image datasets, e.g., GenImage [44], CIFAKE [45], DE-FAKE [30], they focus on fully-synthesized images. Diffusion-based inpainting datasets like COCOGlide [43] and TGIF [46] utilize COCO images but they have limited image quantity, object replacements constrained to the same category (thus lacking out-of-context objects), and no context labels. Our dataset contains significant out-of-context fake objects in real scenes with rich characteristics.
@@ 3 Building COinCO
We propose COinCO, a novel context-oriented inpainted objects dataset derived from the widely-used Common Objects in Context (COCO) Dataset [5]. Leveraging COCO’s image- and object-level annotations, COinCO enhances COCO by systematically replacing objects with diffusion-based inpainting, providing images with enriched contextual diversity. This rendition has 95,320 unique training images and 2,402 testing images from the COCO2017 training and validation sets, respectively. Each inpainted object is categorized as in- or out-of-context and is accompanied by information about the original object and its replacement.
For each COCO image, we (1) randomly pick an object from the image and perform inpainting using Stable Diffusion, (2) verify the new object was successfully inpainted with object detection, and (3) perform context reasoning with an MLLM to classify the new object as in- or out-of-context.
@@ 3.1 Inpainting
(a) Inpainting success rate.
(b) Context reasoning prompt.
Figure 3: (a) Inpainting success rate for original-inpainted object pairs. Rows are classes of original objects and columns are inpainted objects. A darker color indicates higher success rate. (b) Context reasoning prompt for Molmo [17].
We use Stable Diffusion 2 [47] for inpainting. Inpainting a new object requires an original image, an inpainting mask, and a prompt. For each COCO image, we randomly select an object. The object’s mask is slightly dilated and enclosed in a bounding box, which is used as the inpainting mask. In the preliminary study, we observed that the use of dilation and bounding boxes reduced residual artifacts from the original objects. For the replacement object, we randomly select one of COCO’s 80 object categories and use it as the text prompt.
The vanilla Stable Diffusion inpainting pipeline often struggles to inpaint small objects. To overcome this, we crop the original image around an enlarged inpainting area, inpaint, and then scale the results back to the original size. We use alpha blending to seamlessly merge the inpainted region with the original image. For more details on inpainting, see the supplementary materials.
3.2 Automated inpainting verification
Diffusion-based inpainting is imperfect: there is no guarantee the target object will be successfully inpainted. To confirm the success of inpainting, we use a COCO-trained YOLOv8x [48] object detector. If the inpainted object is detected within the inpainting mask, the process is deemed successful. For failed cases, we complete two additional rounds of inpainting and verification, pairing YOLOv8x with a state-of-the-art MLLM (Molmo [17]) to capture YOLO’s false negatives. Images that fail inpainting for all three rounds are discarded, resulting in 97,722 successful images.
Interestingly, we found that inpainting success rate is highly correlated with semantic priors. Figure 3(a) shows the inpainting success proportions for all original-inpainted object pairs. We note dense neighborhoods of inpainted object classes with increased probability of inpainting success. Using average linkage hierarchical clustering, we grouped replacement objects based on inpainting success proportions. With GloVe embeddings, we calculated cosine similarity between items within and across clusters. Mann-Whitney U tests revealed a significantly higher sample mean similarity within clusters (0.169) vs. across clusters (0.119) (bootstrapped p =.$=.002$). The cluster with the highest inpainting success included nearly all animals and food indicating a bias for these classes, while lowest performing cluster included furniture and household objects, which are highly context-dependent. A Kruskal-Wallis test confirmed significant differences across COCO supercategories (p $<0.001$), with food and animals outperforming electronics, appliances, and furniture. These findings validate the semantic coherence of the clusters formed and the vertical block pattern noted in Figure 3(a). Additional statistics are in the supplementary material. To our knowledge, this is the first thorough analysis of semantic priors in diffusion-based inpainting tasks.
@@ 3.3 Context reasoning
We label context for all inpainted objects. An object is considered in-context if it adheres to the contextual coherence of the scene with existing objects, while violations of the coherence result in an out-of-context classification. Our context reasoning is based on three fundamental principles regarding context  [2]: location, size, and co-occurrence. Location examines the spatial positioning of objects within a scene. Size evaluates whether the object’s dimensions are proportionate to the scene geometry. Co-occurrence measures whether objects’ simultaneous presence is common. To label context accordingly, we leverage the power of the Molmo [17]. We design a deliberate prompt to guide Molmo in performing language and visual reasoning for the context classification of our inpainted images (see Figure 3(b)). Provided with an inpainted image and context criteria, Molmo can conduct holistic visual-semantic reasoning that a pure language-based model cannot.
@@ 3.4 Comparison with prior datasets
Table 1: COinCO significantly enriches COCO with inpainting and context labels.
COCO (2017) COCO-Stuff COCOGlide COinCO
@@
Labels objects objects, scenes (fake) objects (fake) objects, context
COinCO extends COCO by introducing contextual diversity with inpainted objects (Table 1). Unlike COCO and COCO-Stuff, which primarily focus on object and scene (stuff) annotations, COinCO provides intentional out-of-context scenarios, enabling research in contextual reasoning and fake localization. While COCOGlide also applies inpainting to COCO, it has significantly less images, and each replacement object has the same class as its original counterpart, thus the context is unaltered.
@@ 4 Manual Verification
To validate the reliability of our automated models for object detection and context classification, we conducted a comprehensive manual verification study. We manually annotated 1,000 images to validate the successful rate of our final dataset, and assessed whether successfully inpainted objects were in- or out-of-context, based on the aforementioned principles.
@@ 4.1 Verification on inpainted object detection
We first evaluated whether target objects can be reliably detected in the inpainted region. We compared COCO-trained detectors, including YOLO [48], MMDetection [49], and GroundingDINO [50] on 1,000 human-labeled images in Table 2. AR: Alignment Rate, PR: Precision, RE: Recall, FPR: False Positive Rate. AR is the primary metric reflecting the agreement between a model and a human.
Method AR (%) PR RE FPR
YOLOv8x 76.4 93.78 67.86 8.15
YOLOv11x 66.6 94.80 50.93 5.06
MMDet 67.6 86.87 58.54 16.01
GroundDINO 59.5 96.14 38.66 2.81
Table 2: Object detection performance.
Model Acc (%) Prec (%) Rec (%) F1 (%)
VisualNet 52.98 51.82 84.62 64.28
SemanticNet 76.33 78.19 73.02 75.52
VisSemanticNet 76.90 77.66 75.51 76.57
Table 3: Context classification performance.
For ground truth, annotators recorded whether the inpainted object was successfully integrated within the mask. We tuned object detector parameters, such as confidence, IoU threshold, and augmentation. YOLOv8x achieved the highest alignment rate (AR) in relation to the manual labels (76.4%)–AR is the rate at which the model agrees with human annotators. While GroundingDINO displayed marginally higher precision (96.14%) and a lower false positive rate (2.81%), the model’s recall was much lower than YOLOv8x (38.66% versus 67.86%), reflecting its failure to identify positives.
In addition, we tested Molmo [17] using cropped inpainting mask regions as input, prompting it to confirm whether the target object was present. Molmo achieved an AR of $82.6\%$ with human annotations, surpassing the performance of conventional object detectors. Thus, YOLOv8x and Molmo work together in our cascaded verification process for efficiency and accuracy.

### v2

@@ Abstract
We present Common Inpainted Objects In-N-Out of Context (COinCO), a novel dataset addressing the scarcity of out-of-context examples in existing vision datasets. By systematically replacing objects in COCO images through diffusion-based inpainting, we create 97,722 unique images featuring both contextually coherent and inconsistent scenes, enabling effective context learning. Each inpainted object is meticulously verified and categorized as in- or out-of-context through Large Vision Language Model assessments. We demonstrate three key tasks enabled by COinCO: (1) a fine-grained context reasoning approach that classifies objects as in- or out-of-context based on three criteria; (2) a novel Objects-from-Context prediction task that determines which new objects naturally belong in given scenes at both instance and clique level semantics, and (3) context-enhanced fake detection on state-of-the-art methods without fine-tuning. COinCO provides a controlled testbed with contextual variations, establishing a foundation for advancing context-aware visual understanding in computer vision, including image forensics. Code and dataset are available at https://co-in-co.github.io/.
@@ 1 Introduction
Context is fundamental to visual understanding [3, 43, 45, 66, 13, 61]. When humans view a scene, we instinctively assess the coherence between objects and their environment. This context-based reasoning is essential for judging whether objects naturally belong in a scene [5]. Take a look at Figure 1—do the red-boxed objects seem reasonable in their contexts? Context reasoning can reveal out-of-place objects, providing a complementary cue for detecting manipulated content. By using context, humans assess whether objects appear in plausible settings—a horse in a grassy field with trees in the background appears perfectly natural, while a tiny zebra on a beach next to a person immediately raises suspicion due to its unusual size and location.
@@
Learning context from data, however, is difficult. A significant challenge is that unusual scenes with out-of-context objects are rare in real life. Common computer vision datasets [34, 10, 21, 31] primarily contain objects in their natural settings, resulting in a scarcity of examples with contextual violations. This presents a fundamental obstacle: how can we train machine learning models that require large data to recognize contextual inconsistencies when such examples are inherently uncommon?
Figure 1: COinCO contains a rich set of inpainted out-of-context objects (top two rows) and in-context objects (bottom row).
@@
To address this challenge, we propose a new dataset with both in-context and out-of-context objects through controlled image manipulation. By systematically replacing objects in existing real scenes, we generate the necessary examples of contextual violations while preserving the overall structure of the scene. We specifically build on top of the Common Objects in Context (COCO) dataset [34] because it contains a diverse set of everyday scene photographs with rich image- and object-level annotations, offering an ideal starting point for contextual manipulation.
@@
Using a diffusion-based inpainting model [48], we replace exactly one object per COCO image. This selective approach allows us to maintain the broader scene context while introducing precise, controlled variations in object-scene relationships. Our meticulous data creation pipeline and multi-step verification ensure high-quality inpainting results, and we apply a state-of-the-art Large Vision Language Model (LVLM) to classify each inpainted object as either contextually consistent or inconsistent with its scene.
@@
We name our dataset Common Inpainted Objects In-N-Out of Context (COinCO), emphasis on context for inpainted objects. It contains 97,722 inpainted images, each accompanied by rich annotations. Figure 2 shows the pipeline of our data creation and the downstream tasks.
@@
COinCO is the first dataset that features inpainted objects with diverse context annotations, enabling advancements in tasks including fine-grained context classification, object-from-context prediction, and fake detection (as our objects are inpainted by a generative model).
Our main contributions are:
@@
2. We introduce a fine-grained context reasoning task in COinCO based on three criterion-specific dimensions and train efficient student models through distillation from a 72B LVLM teacher. These models provide modular, interpretable, and high-accuracy context reasoning for inpainted and real images.
3. We introduce a novel task, object-from-context, which aims to predict instance- and clique-level semantic categories for possible objects that fit the given context.
4. We show that context can be effectively integrated into fake detection pipelines, substantially improving state-of-the-art image forensics methods without fine-tuning.
Figure 2: COinCO pipeline. Left: Building COinCO. We start with COCO images and annotations. For each image, we randomly select an object and replace it using Stable Diffusion inpainting. We verify the inpainted object using object detection and a vision-language model. Successfully inpainted images are added to the dataset, while failed cases are regenerated and retested. Finally, each inpainted object is labeled as in-context or out-of-context based on three criteria: location, size, and co-occurrence. Right: COinCO Tasks. We demonstrate three downstream applications enabled by COinCO. Sec 5.1: Fine-grained context classification evaluates objects on each criterion using vision-language models. Sec 5.2: Objects-from-context prediction identifies which object categories fit a given context at both instance and clique level semantics, and can recover the original object with an image generation model. Sec 5.3: Context-enhanced fake detection improves pretrained fake detectors by enhancing predictions in out-of-context regions.
@@ 2 Related Work
Context reasoning. Context is essential for understanding object relationships and finding anomalies in complex scenes [43]. Biederman et al. [5] identified relational principles such as support, probability, and size that help identify contextual inconsistencies, which prior works apply using co-occurrence and support relationships for out-of-context detection [13]. Others have shown that slight context changes can cause significant errors in object detection, and Acharya et al. [1] introduced a Graph Contextual Reasoning Network modeling co-occurrence and relative position to detect out-of-context objects. Context is particularly important for object detection [17, 7]; a comprehensive review is in [58]. Ours is the first work to feature a mixture of in- and out-of-context objects in large-scale data. While context remains under-explored in fake detection, Large Vision Language Models (LVLMs) [15, 36, 35, 54, 28] offer a human-aligned approach to context reasoning; we are the first to leverage them for context-based fake detection.
Fake image generation and detection. Image manipulation has evolved from traditional object insertion [18, 9] to GAN-based synthesis [19, 25, 8, 29, 30] and diffusion models [23, 49], including Stable Diffusion [48], DALL-E [46], and ControlNet [65]. Emerging techniques further tailor diffusion methods for object manipulation, inpainting, and harmonization [11, 64, 52], 3D [37], and relighting [22, 26]. For detection, earlier works classify at image-level [14, 42], while advanced methods localize manipulations: PSCC-Net [39] uses spatio-channel correlation, CAT-Net [32] exploits JPEG artifacts, ManTra-Net [60] captures manipulation traces, and TruFor [20] combines RGB and noise-sensitive fingerprints. For fake image datasets, e.g., GenImage [67], CIFAKE [6], DE-FAKE [49], they focus on fully-synthesized images. Diffusion-based inpainting datasets like COCOGlide [20] and TGIF [41] utilize COCO images but they have limited image quantity, object replacements constrained to the same category (thus lacking out-of-context objects), and no context labels. Our dataset contains significant out-of-context fake objects in real scenes with rich characteristics.
@@ 3 Building COinCO
We propose COinCO, a novel context-oriented inpainted objects dataset derived from the widely-used Common Objects in Context (COCO) Dataset [34]. We specifically choose COCO as our foundation for several reasons. First, we focus on common, everyday objects where contextual reasoning is most meaningful and widely applicable. COCO’s 80 classes span diverse supercategories covering frequent real-world categories such as animals, food, vehicles, and household items, making it the standard benchmark for context-rich scenes. Second, COCO provides reliable instance masks and annotations essential for high-quality inpainting. Expanding to rare or unusual objects would dilute the contextual focus and introduce annotation noise, as contextual expectations become less defined for uncommon categories. Leveraging COCO’s image- and object-level annotations, COinCO enhances COCO by systematically replacing objects with diffusion-based inpainting, providing images with enriched contextual diversity. Our dataset comprises 97,722 successfully inpainted images derived from COCO2017, with each inpainted object categorized as in- or out-of-context with reasoning and accompanied by information about the original object and its replacement. Task-specific evaluation protocols are detailed in their respective sections.
For each COCO image, we (1) randomly pick an object from the image and inpaint a new object, (2) verify the new object was successfully inpainted with object detection, and (3) perform context reasoning with LVLMs to classify the new object as in- or out-of-context.
@@ 3.1 Inpainting
We use Stable Diffusion [48] for inpainting. Inpainting a new object requires an original image, an inpainting mask, and a prompt. For each COCO image, we randomly select an object. The object’s mask is slightly dilated and enclosed in a bounding box as the inpainting mask. Dilation and bounding boxes reduce residual artifacts from the original objects. For replacement, we randomly select one of COCO’s 80 categories as the text prompt.
The vanilla inpainting pipeline often struggles to inpaint small objects. To overcome this, we crop the original image around an enlarged inpainting area, inpaint, and then scale the results back to the original size. We use alpha blending to seamlessly merge the inpainted region with the original image. We found that for stronger inpainting models, such as Flux [33] and Qwen Image [59], they focus more on visual fidelity on in-context objects and tend to sacrifice the ability to generate out-of-context objects and provide less precise spatial control over the inpainting region. More details on inpainting and the analysis of different inpainting models are in the supplementary materials.
Figure 3: Context reasoning prompt for LVLMs.
@@ 3.2 Context reasoning
We annotate context labels to all inpainted objects through a multi-model consensus process. Rather than relying on one single large model, three state-of-the-art LVLMs (Molmo-72B [15], Qwen2.5-VL-72B [2], and InternVL3.5-38B [57]) independently evaluate each inpainted image according to three context criteria: location, size, and co-occurrence. We design a structured prompt (Figure 3) to guide the LVLMs in performing this contextual reasoning. For every image, each model first provides criterion-level reasoning by examining (1) whether the spatial placement of the object is appropriate for the scene layout, (2) whether its scale aligns with the scene geometry, and (3) whether its simultaneous presence with existing objects is contextually plausible. An object is labeled as out-of-context if it violates any of these criteria; otherwise, it is deemed in-context.
After generating criterion-level analyses, each LVLM produces a final binary judgment (in-context vs. out-of-context). To ensure the reliability of our context annotations, we retain labels only for images where a unanimous agreement is reached. This consensus requirement yields 73,929 images for which we release both the final in-/out-of-context labels and the complete criterion-level reasoning from all LVLMs. The remaining images are still part of COinCO, but we do not provide context annotations for them due to labeling ambiguity.
@@ 3.3 Comparison with prior datasets
Table 1: COinCO significantly enhances COCO in new ways.
COCO (2017) COCO-Stuff COCOGlide Ours
@@
Labels objects objects, scenes (fake) objects (fake) objects, context reasoning
COinCO extends COCO by introducing contextual diversity with inpainted objects (Table 1). Unlike COCO and COCO-Stuff, which primarily focus on object and scene (stuff) annotations, COinCO provides intentional out-of-context scenarios, enabling research in contextual reasoning and fake localization. While COCOGlide [20] also applies inpainting to COCO, it has significantly less images, and each replacement object has the same class as its original counterpart, thus the context is unaltered.
@@ 4 Manual Verification
To validate the reliability of our automated models for object detection and context classification, we conducted a comprehensive manual verification study. We manually annotated 1,000 images to validate the successful rate of our final dataset, and assessed whether successfully inpainted objects were in- or out-of-context from human judgment.
@@ 4.1 Verification on inpainted object detection
To select an appropriate detector for our automated inpainting verification pipeline, we evaluated several COCO-trained models on 1,000 human-labeled images where annotators recorded whether the target object was successfully inpainted within the mask region. As shown in Table 2, while GroundingDINO [38] achieved the highest precision (96.14%), its low recall (38.66%) would discard too many successful inpainting results. YOLOv8x [47] achieved the best balance with an F1 score of 78.75%, combining high precision (93.78%) with reasonable recall (67.86%). We therefore selected YOLOv8x as our detector and use it to verify inpainted objects: if YOLOv8x detects the target object within the inpainting mask, the process is deemed successful; for failed cases, we complete up to two additional rounds of inpainting and verification; images that fail all three rounds are discarded, resulting in 97,722 successful images. Notably, false positives often correspond to lower-quality inpaintings where the target object is present but poorly rendered, rather than complete inpainting failures.
Table 2: Object detection performance on inpainted objects.
Metric YOLOv8x YOLOv11x MMDet GroundDINO
Precision 93.78 94.80 86.87 96.14
Recall 67.86 50.93 58.54 38.66
F1 78.75 66.26 69.95 55.14